OpenStudy (summithill):

Joe drives 20mi in 1/2h how long will it take him to drive 50mi?

OpenStudy (jackthegreatest):

50/20 = ?

OpenStudy (summithill):

hmm

OpenStudy (summithill):

2.5

OpenStudy (jackthegreatest):

1/2 hr = how many min

OpenStudy (summithill):

30

OpenStudy (jackthegreatest):

30 x 2.5 = ?

OpenStudy (summithill):

75 minutes im guessing

OpenStudy (jackthegreatest):

yes

OpenStudy (summithill):

simplify right

OpenStudy (jackthegreatest):

yes to hours

OpenStudy (jackthegreatest):

and your done

OpenStudy (summithill):

1 hour and 15 minutes ;D

OpenStudy (summithill):

You get a medal ;D

OpenStudy (jackthegreatest):

correct

OpenStudy (triciaal):

alternate approach to provide explanation and understand the process here you need to set up a ratio

OpenStudy (summithill):

ok

OpenStudy (jackthegreatest):

20/30 = 50/x @triciaal :D stole your idea

OpenStudy (summithill):

Lol

OpenStudy (triciaal):